Mary's age is 5 less than twice the age of her younger sister, Grace. The sum of their ages is 19. What is the age of each?

Answer by satyareddy22(84)

Let the ages of mary and grace be x and y respectively.
Mary's age is 5 less than twice the age of her younger sister
2y-x=5-----(1)
sum of their ages is 19
y+x=19-----(2)
From (1)x=2y-5,substituting in (2)
y+2y-5=19
3y=19+5=24
3y=24=>y=24/3=8
x=2y-5=2*8-5=16-5=11
So the ages of mary and grace are 11yrs and 8yrs respectively